A child sits at a distance of 5.0 m from the axis of merry-go-round in an amusement park. If the merry-go-round makes 2 rev



in 10.0s, find the (a) tangential speed and (b) centripetal acceleration of the child.

1
Expert's answer
2022-04-06T18:36:44-0400

(a) tangential speed

"v=\frac{2\pi R}{T}=\frac{2\pi*5.0}{5.0}=6.28\:\rm m/s"

(b) centripetal acceleration

"a=\frac{v^2}{R}=\frac{6.28^2}{5.0}=7.9\:\rm m/s^2"